Reports 1-1 of 1 Clear search Modify search
AdV-INJ (Input Mode Cleaner cavity)
spinicelli, gosselin, chiummo - 21:02 Friday 11 September 2020 (49402) Print this report
IMC AA sensing matrix measurement

After the realignment of yesterday (49393) and the improvement of the PMC analog input, PMC was quite stable overnight (fig. 1). Nevertheless, it was impossible to close the AA loop in full bandwidth.

We tried then to recover IB and MC positions to which of the last AA matrix sensing of Monday 7, but correction signals were far from 0, in particular IB_Tx. Moreover, IMC transmission (IMC_TRA) was lower than what we expected from imput power (~24.3 W vs ~25.3 W) . We could get back the transmitted power and a good reflection in IMC_ref camera by realigning both IB and MC,  even though  we only partially improved correction signals. Then, we manually tuned the Fmoderr and we measured again a new sensing matrix for AA following the procedure in 49366:

Injection time
MC ty gps = 1283867415.00
IB ty gps = 1283866846.00
MC tx gps = 1283867329.00
IB tx gps = 1283866773.00
IB tz gps = 1283866968.00

With the new matrix, we could close the IMC Aa but not in full bandwidth yet because IB_tx, the control with higher correction signal, started to oscillate. We decided to leave IMC locked with drift control, closing RFC and PSTAB for the weekend (fig. 2).

We will need to investigate it on Monday.

Images attached to this report
Search Help